Paul is 12. His father is 33. How many years ago was the father 4 times as old as the son?

Answer by ewatrrr(24785)


Hi
Paul is 12. His father is 33
Let x represent How many years ago was the father 4 times as old as the son?
Question states*** 'x' years ago...
33-x = 4*(12-x)
Solving for x
33-x = 48 - 4x
3x = 15
x = 5, the years ago the father was 4 times as old as the son
CHECKING our Answer***
33-5 = 28 = 4*(12-5)